### v1

# Determination of the resonant parameters of $\chi_{c0}(3915)$ with global fit

Abstract
The $\chi_{c0}(3915)$ was observed firstly by the Belle experiment in the $\omega J/\psi$ invariant mass spectrum in the process $B\to K\omega J/\psi$, and then confirmed by the Babar experiment. The two experiments reported the resonant parameters of this particle in the both processes $\gamma\gamma\to\omega J/\psi$ and $B\to K\omega J/\psi$ by assuming $\chi_{c0}(3915)$ as a S-wave Breit-Wigner resonance. We perform a global fit to the distributions of invariant mass of $\omega J/\psi$ measured by the Belle and Babar experiments, and incorporate the measurement by the LHCb experiment additionally to extract the mass and width of $\chi_{c0}(3915)$. We obtain $M=3921.0\pm 0.9$ MeV/$c^{2}$ and $\Gamma=17.9\pm 2.2$ MeV which are in good agreement with the values on PDG but with higher precision.
I Introduction
Since the first charmonium-like state $X(3872)$ was discovered by Belle experiment via the process $e^{+}e^{-}\to\gamma_{ISR}\pi^{+}\pi^{-}J/\psi$, a new era of the study of charmonium-like states was triggered [1]. A series of mesons comprised of charmed and anticharmed quark pairs, such as the Y(4260), $X(3915)$, and Zc(3900), were subsequently discovered, and most of them were confirmed by different experiments [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The understanding of the nature of these particles has turned out to be quite a challenge. Many of them have properties that are quite different from the conventional charmonium, e.g. low open-charm decay rate or absence of hadronic transitions to other charmonium states. To explain these anomalous features, many models have been proposed by theorists, including charmonium molecule mixed states [12, 13], $c\bar{c}g$ hybrid states [15], and tetra-quarks [14]. In addition, the production rates for some of these charmonium-like states are quite low in most experiments which limits the measurement precision of the resonant parameters and the determination of the corresponding quantum numbers, which make the interpretation of these particles difficult. Particle Data Group (PDG) [16] has assigned the $X(3915)$ and $X(3872)$ as the excited spin-triplet charmonium states $\chi_{c0}(2P)$ and $\chi_{c1}(2P)$ named $\chi_{c0}(3915)$ and $\chi_{c1}(3915)$, respectively. We just follow the PDG name convention in the article.
As a member of charmonium-like family of states, the $\chi_{c0}(3915)$ was first observed by the Belle experiment in the process $B\to KJ/\psi\omega$ in a data sample containing $275\times 10^{6}$ $B\bar{B}$ pairs [9]. The measured mass and width were determined to be $3943\pm 11\pm 13$ MeV/$c^{2}$ and $87\pm 22\pm 26$ MeV with the assumption that the $\chi_{c0}(3915)$ is an S-wave Breit-Wigner (BW) resonance. The particle was confirmed by Babar experiment in the same decy mode with a $383\times 10^{6}$ $B\bar{B}$ event data sample [10], their reported mass and width were $3914.6^{+3.8}_{-3.4}\pm 2.0$ MeV/$c^{2}$ and $34^{+12}_{-8}\pm 5$ MeV. Babar experiment subsequently updated their measurements with a lager data sample of $467\times 10^{6}$ $B\bar{B}$ events and looser $M(\pi^{+}\pi^{-}\pi^{0})$ requirements that revealed a X(3872) signal as well [11]; the updated mass and width were $3919.1^{+3.8}_{-3.4}\pm 2.0$ MeV/$c^{2}$ and $31^{+10}_{-8}\pm 5$ MeV. In addition, the $\chi_{c0}(3915)$ was also observed in the two-photon collision process $\gamma\gamma\to\omega J/\psi$ by both the Babar and Belle experiments [7, 8]. Their measured masses and widths are listed in Table 1. Babar did a spin-parity analysis in their analysis and obtained the quantum number to be $J^{P}=0^{+}$, and identified the $\chi_{c0}(3915)$ as the $\chi_{c0}(2P)$ charmonium state. However, this assignment was disputed because of the large rate for $\chi_{c0}(3915)\to\omega J/\psi$ decay and the absence of $\chi_{c0}(3915)\to D\bar{D}$ decays [17, 18]. Moreover, the mass difference between the $\chi_{c2}(2P)$ and $\chi_{c0}(3915)$ is only about 10 MeV, which is too small for the fine splitting of $P$-wave charmonia [17]. In 2020, LHCb experiment made an amplitude analysis of the $B^{+}\to D^{+}D^{-}K^{+}$ decay [19], and reported that the spin-0 resonance is of necessity to describe the data well and determined its mass and width to be $2923.8\pm 1.5\pm 0.4$ MeV/$c^{2}$ and $17.4\pm 5.1\pm 0.8$ MeV, respectively.
In this article, we perform a simultaneous $\chi^{2}$ fit to the distributions of invariant mass of $\omega J/\psi$ in the processes of $\gamma\gamma\to\omega J/\psi$ measured by Belle [denoted as (a)], $\gamma\gamma\to\omega J/\psi$ by Babar [(b)], $B^{0}\to\omega J/\psi K^{0}$ by Babar [(c)], $B^{+}\to\omega J/\psi K^{+}$[(d)] by Babar, and $B\to\omega J/\psi K$ by Belle [(e)] to extract the mass and width of $\chi_{c0}(3915)$. The distributions of $M(\omega J/\psi)$ for these processes are shown in Figure 1. Furthermore, LHCb’s results are taken into account as an additional component in the $\chi^{2}$ calculation. Compared to the values on PDG who also gives the $\chi_{c0}(3915)$ mass and width by combining the measurements by these experiments, we use more detailed information of the $\omega J/\psi$ invariant mass spectrum which expect to provide a result of higher precision.
Figure 1: Fit to the distributions of the $\omega J/\psi$ invariant mass. The black dots with error bars are the data mentioned in the main text. The red solid curves are the fit results, the cyan/blue dashed curves indicate the $\chi_{c1}(3872)$/$\chi_{c0}(3915)$ signals, and the green dashed curves show the backgrounds.
II simultaneous $\chi^{2}$ fit
In the simultaneous $\chi^{2}$ fit, the functions used to fit the $M(\omega J/\psi)$ spectra in Fig. 1 are similar to those applied in their corresponding publications. The fit function is comprised of signal and background components. The $\chi_{c0}(3915)$ signal shape is described by a S-wave BW function convoluted with the detector resolutions. The BW is $\frac{\Gamma(p^{*}/p_{0})}{(m^{2}-M^{2})^{2}+(M\Gamma(p^{*}/p_{0}))^{2}}$, where $M$ is the peak mass, $p^{*}$ is the momentum of $J/\psi$ momentum in the rest frame of a $\omega J/\psi$ system, and $p_{0}=p^{*}$ when $m=M$ [20]. $M$ and $\Gamma$ are common parameters for the four measurements that are allowed float in the fitting. The contributions of the $\chi_{c1}(3872)$ are needed for the measurements (c) and (d) as shown in Fig. 1. The signal shape of the $\chi_{c1}(3872)$ is described with a simple BW function with a fixed mass and width of $3871.69$ MeV/$c^{2}$ and 1.2 MeV, respectively. The detector resolution of $\gamma\gamma\to\omega J/\psi$ for Belle is described by a double Gaussian function, one Gaussian has the mean and deviation values of 4.5 MeV and 0 MeV, respectively, with a coefficient of 0.59, and the other one has 16 MeV and -4.0 MeV with a cofficient of 0.41 [7]. The resolution for the process $\gamma\gamma\to\omega J/\psi$ by Babar is described by a single Gaussian with a deviation of $6.7$ MeV and a mean value of zero. The resolution for the measurement $B\to K\omega J/\psi$ by Babar is described by a single Gaussian function with a deviation of 5.7 MeV and a mean value of zero. All of the resolution information is from the corresponding publications. The non-resonance background shapes for the measurement (a) and (b) are both described as $p^{*}(m){\rm exp}(-\delta p^{*}(m))$, where $\delta$ is a parameter that is allowed to float in the fitting, and $m=m(J/\psi\omega)$ [8]. The background shapes of the measurement (c) and (d) are described with a Gaussian function with parameters that float in the fitting [11]. The shape of the background in the measurement (e) is described with the $J/\psi$ momentum in the rest frame of a $\omega J/\psi$ system, $p^{*}(M)$. The mass and width are the common parameters in all the $\chi_{c0}(3915)$ BW functions and allowed to be free in the fit.
The $\chi^{2}$ is defined as
$\chi^{2}=\sum_{i=1}^{5}\sum_{j=1}^{N_{i}}(\frac{x_{ij}-\mu_{ij}}{\sigma_{ij}})^{2}+(\frac{M_{h}-M}{\sigma_{M_{h}}})^{2}+(\frac{\Gamma_{h}-\Gamma}{\sigma_{\Gamma_{h}}})^{2}$ (1)
where $i$ have values from 1 to 5, corrsponding to the measurements (a) to (e), $x_{ij}$ and $\sigma_{ij}$ are the observed signals and corresponding errors in each bin as shown in Fig. 1, $\mu_{ij}$ is the expected value in each bin calcluated with the fitting shape mentioned above, and $N_{i}$ is the number of bins in each measurement. The bin width of the measurements (a) and (b), shown in the published papers [7, 8], is 10 MeV, however, the number of observed events in each bin in the mass region of 3.95 GeV to 4.3 GeV is very small, so the bin width is reset to be 50 MeV in this region to ensure that statistics in each bin are large enough to make a meaningful calculation of their contribution to the $\chi^{2}$ value. The numbers of bins in the measurement (a) and (b) are $N_{1}$=19 and $N_{2}$=15. For the measurements (c) and (d), the bin width in the mass region of 3.8425 to 3.9925 GeV/$c^{2}$ is 10 MeV/$c^{2}$, and 50 MeV/$c^{2}$ in the region beyond 3.9925 GeV/$c^{2}$ as shown in Fig. 1. The bin numbers $N_{3}$ and $N_{4}$ are both 31. For the measurement (e), $N_{5}$=12. The rest two components of the $\chi^{2}$ formula, $\frac{M_{h}-M}{\sigma_{M_{h}}}$ and $\frac{\Gamma_{h}-\Gamma}{\sigma_{\Gamma_{h}}}$, are the measurements from the LHCb experiment, where $M_{h}$ and $\Gamma_{h}$ are the measured mass and width, and $\sigma_{M_{h}}$ and $\sigma_{\Gamma_{h}}$ the corresponding statistical uncertainties.
By minimizing the $\chi^{2}$ with minuit [21], we obtained the fit results with $M=3921.0\pm 0.8$ MeV/$c^{2}$ and $\Gamma=17.9\pm 2.1$ MeV. The goodness of fit is $\chi^{2}/ndf=113.1/87$ where $ndf$ is the number of degree of freedom in the fit. The fit results are also shown in Fig. 1.
The total systematic uncertainty of the mass and width, $\sigma_{\rm sum}$, is obtained with the formula $\frac{1}{\sigma_{\rm sum}^{2}}=\sum_{i=1}^{5}\frac{1}{\sigma_{i}^{2}}$, where the $i$ has a value from 1 to 5 corresponding to each measurement listed in Table 1, and $\sigma_{i}$ is the systematic uncertainty of the $i$-th measurement. Combining with the statistical uncertainties from the previous fit, we finally get the results of $M=3921.0\pm 0.9$ MeV/$c^{2}$ and $\Gamma=17.9\pm 2.2$ MeV.
Table 1: Mass ($M$) and width ($\Gamma$) of the $\chi_{c0}(3915)$ measured by the different experiments, where the first uncertainty is the statistical and the second is systematic. The results from PDG and this work are listed in the last two rows.
Experiment Production $M$ (MeV/$c^{2}$) $\Gamma$ (MeV)
Belle [9] $B\to\omega J/\psi K$ $3943\pm 11\pm 13$ $87\pm 22\pm 26$
Babar [11] $B\to\omega J/\psi K$ $3919.1^{+3.8}_{-3.4}\pm 2.0$ $31^{+10}_{-8}\pm 5$
Belle [7] $\gamma\gamma\to\omega J/\psi$ $3915\pm 3\pm 2$ $13\pm 6\pm 3$
Babar [8] $\gamma\gamma\to\omega J/\psi$ $3919.4\pm 2.2\pm 1.6$ $17\pm 10\pm 3$
LHCb [19] $B^{+}\to D^{+}D^{-}K^{+}$ $2923.8\pm 1.5\pm 0.4$ $17.4\pm 5.1\pm 0.8$
PDG [16] $3921.7\pm 1.8$ $18.8\pm 3.5$
This work $3921.0\pm 0.9$ $17.9\pm 2.2$
III summary
In summary, we determine the resonant parameters of $\chi_{c0}(3915)$ by simultaneously fitting the measurements provided by the Belle, Babar, and LHCb experiments. The mass and width are determined to be $M=3921.0\pm 0.9$ MeV/$c^{2}$ and $\Gamma=17.9\pm 2.2$ MeV, respectively, which are in good agreement the average values in PDG but with higher precision.
Acknowledgments
This work is supported by National Natural Science Foundation of China (NSFC) under contract No. 11805090.